OpenStudy (anonymous):

Find the x-intercept of y = 4x - 64. A) (3, 0) B) (64, 0) C) (0.602, 0) D) (1.806, 0)

OpenStudy (anonymous):

to find x intercept, set y=0 and solve for x to find y intercept, set x=0 and solve for y

OpenStudy (anonymous):

******************************** y = 4x - 64 ********************************* To find the x-intercept, let y = 0 y = 4x - 64 0 = 4x - 64 ********************************** To find the y-intercept, let x = 0 y = 4x - 64 y = 4(0) - 64 ***********************************

OpenStudy (anonymous):

have you figured it out yet @KendraSteed123 ?